# A new machine learning framework for occupational accidents forecasting with safety inspections integration

### v1

@@ Abstract
Introduction: Reducing the number of occupational accidents remains a major challenge for companies, as these events lead to significant human harm and financial losses. Although many organizations have implemented safety programs and made continuous efforts to improve their prevention strategies, these measures often remain insufficient to proactively and dynamically anticipate risks. In particular, safety inspections are still largely underexploited, and their integration into continuously updated predictive models has received little attention. Methods: we propose a generic framework for short-term occupational accident forecasting that leverages safety inspections and models accident occurrences as binary time series. The approach generates daily predictions, which are then aggregated into weekly safety assessments for better decision making. To ensure the reliability and operational applicability of the forecasts, we apply a sliding-window cross-validation procedure specifically designed for time series data, combined with an evaluation based on aggregated period-level metrics. Several machine learning algorithms, including logistic regression, tree-based models, and neural networks, are trained and systematically compared within this framework. Results:the long short-term memory (LSTM) network outperforms the other approaches and detects the upcoming high-risk periods with a balanced accuracy of 87%, confirming the robustness of our methodology and demonstrating that a binary time series model can anticipate these critical periods based on safety inspections. Conclusions and Practical Applications: the proposed methodology converts routine safety inspection data into clear weekly risk scores, detecting the periods when accidents are most likely to occur.Decision-makers can integrate these scores into their planning tools to classify inspection priorities, schedule targeted interventions, and funnel resources to the sites or shifts classified as highest risk, stepping in before incidents occur and getting the greatest return on safety investments.
@@ 1 Introduction
The international labour organization (ILO) estimates that nearly 300,000 people die each year due to occupational accidents (ILO, 2023). In France, the National Health Insurance recorded over 600 000 occupational accidents in 2023, nearly 700 of them fatal and this level has remained essentially flat for over a decade (Amelie, 2024). This observation highlights the limitations of current strategies and the urgent need for new approaches to sustainably reduce both the frequency and severity of occupational accidents.
@@
Since Heinrich’s pioneering work (Heinrich, 1931) and his domino theory, the understanding of occupational accidents has evolved considerably. These events are no longer seen as isolated or random, but rather as the outcome of a chain of contributing factors with complex interactions. Several theoretical models have emerged, such as Surry’s sequence of events model (Surry, 1969), Reason’s Swiss cheese model (Reason, 1990), and the cause tree method developed by the French INRS institute (INRS, 2019). While these models have helped structure accident investigation processes, they remain primarily retrospective, lack predictive capability, and often fail to capture the temporal dynamics of risk in complex environments (Qureshi, 2007; Larouzee and Le Coze, 2020).
@@
With the growing adoption of machine learning, new proactive strategies have been introduced in sectors such as construction, mining, agriculture, and services. Numerous studies demonstrate the potential of these techniques for incident prediction (Suárez Sánchez et al., 2011; Rivas et al., 2011; Wang et al., 2019), risk assessment (Palei and Das, 2009; Weng and Meng, 2011; Leu and Chang, 2013), injury severity classification (Chang and Chien, 2013; Esmaeili et al., 2015; Tixier et al., 2016), and risk-factor identification (Cheng et al., 2012; Amiri et al., 2016). However, most of these approaches rely on lagging indicator data collected after an accident, which limits their usefulness for anticipating risks and taking preventive action in advance. To address this limitation, the concept of safety leading indicators (SLIs) has gained momentum. In contrast to lagging indicators, SLIs allow for the early detection of weak signals based on proactive field data (Reiman and Pietikäinen, 2012; Hinze et al., 2013). Numerous studies emphasize their usefulness for constructing predictive models (Grabowski et al., 2007; Poh et al., 2018; Gondia et al., 2023). However, most of the existing frameworks using SLIs suffer from two major limitations: (1) they are rarely updated continuously, limiting their ability to adapt to evolving operational contexts; and (2) they often fail to explicitly capture temporal dependencies and to integrate recent information into the prediction process. Time series models have also been applied to accident forecasting (Carnero and Pedregal, 2010; Koc et al., 2022), but they typically operate at national or regional levels, over long time spans. Other works focus on building early warning systems based on composite indicators (Li et al., 2016; Nazaripour et al., 2018), yet these systems require domain-expert thresholds and domain-specific calibration, limiting their operational flexibility. In this paper, we introduce a generic framework designed to enable short-term forecasting of occupational accidents at a company level, both globally and by its organizational units. The framework models daily accident occurrences as a binary time series (Kedem and Fokianos, 2002; Fokianos and Kedem, 2003), where each day is represented by a binary outcome (accident or no accident), predicted using machine learning algorithms. Forecasts are produced on a daily basis and subsequently aggregated to provide a weekly risk assessment using metrics adapted to imbalanced classification problems (Luque et al., 2019). A sliding-window cross-validation procedure, tailored for time series data (Tashman, 2000; Bergmeir and Benítez, 2012; Hyndman and Athanasopoulos, 2018), is employed to assess the model robustness and practical applicability in operational safety monitoring. The proposed framework enables dynamic weekly assessments of whether a given upcoming period is likely to be safe or risky, based on real-time data collected from the field. To determine the weekly risk status, the system estimates daily probabilities of accident occurrence. These are then binarized using a calibrated threshold: a day is flagged as at risk if its predicted probability exceeds this threshold, and a week is classified as risky if it includes at least one such day.
For illustration, Table 1 shows the output of our approach applied to a specific department within the company under study, over a two-week period. The daily threshold for binarizing accident risk is set based on a calibration process and fixed here at $0.6$. In week 3, two days exceed this threshold, resulting in the classification of the entire week as risky. In contrast, in week 4, no day probability crosses the threshold, so the week is classified as safe. This example highlights how the approach can be used at the departmental level to provide timely and actionable insights for occupational risk management.
Week 3 Week 4
Days of the week M T W T F S S M T W T F S S
Daily probability 0.996 0.620 0.002 0.001 0.012 0.001 0.013 0.000 0.001 0.000 0.000 0.000 0.000 0.004
Forecast 1 1 0 0 0 0 0 0 0 0 0 0 0 0
@@
Table 1: Output of the proposed approach for a single department over two consecutive weeks. Weeks are flagged as risky or safe by comparing each day’s accident probability with a calibrated threshold of 0.6; week 3 is labelled risky because two days exceed the threshold, whereas week 4 remains safe since no day does.
@@ 2 Literature review
An indicator is a qualitative or quantitative measure used to assess or monitor the evolution of a situation, phenomenon, or activity. In the field of data-driven occupational safety management, two main types of indicators are typically distinguished (Grabowski et al., 2007; Hopkins, 2009): lagging indicators and leading indicators. Lagging indicators such as accident or incident rates, compensation costs and number of injuries resulting in time off work (Choudhry et al., 2007; Hinze et al., 2013; Jazayeri and Dadi, 2017) reflect the consequences of accidents that have already occurred. In other words, they are updated only after an accident happens. Several authors (Grabowski et al., 2007; Mengolini and Debarberis, 2008) argue that such indicators do not provide sufficiently useful information to prevent future accidents. According to Lindsay (1992), a low number of reported accidents even over several years does not necessarily mean that risks are under control or that other incidents will not occur. Despite their limitations, these indicators remain widely used because they are easy to quantify and identify (Lingard, 2013; Almost et al., 2018) and allow organizations to benchmark against one another. (Elsebaei et al., 2020).
In contrast, leading indicators provide early warning signs of accidents and adopt a more proactive approach, aiming to detect and act before incidents occur (Mearns, 2009; Eaton et al., 2013). Examples include near-miss report, safety talks, and safety inspections (Falahati et al., 2020). Field-level feedbacks can also be added to this list, as they enable quick and spontaneous collection of real-world operational data, helping to manage weak signals in real time. In the construction sector, Hinze et al. (2013) highlight the importance of these indicators and differentiates between passive leading indicators (e.g., number of employees trained or presence of a prevention plan) and active leading indicators (e.g., the percentage of safety meetings attended by supervisors). The latter reflects more accurately the dynamic reality of prevention efforts.
@@
Despite their promise, leading indicators remain difficult to adopt widely, partly due to the diversity of work environments: an effective indicator in construction may not apply in agriculture or maritime industries (Hinze et al., 2013; Xu et al., 2021). Furthermore, the subjectivity associated with some indicators such as the assessment of the severity of a hazardous situation can distort the actual perception of system performance or activity (Grabowski et al., 2007). Sometimes, the boundary between leading and lagging indicators is blurry, and some indicators are poorly defined or misaligned with their intended objectives. It is therefore crucial to distinguish between process safety hazards risks inherent to the operation of the system (e.g., explosions or toxic spills) and personal safety hazards which are more related to individual accidents such as falls, crushes, or electrocutions (Hopkins, 2009). Common lagging indicators like accident rates are often focused on personal safety and fail to capture process-related risks effectively. Similarly, some leading indicators (e.g., audit frequency) can remain too generic if they do not account for the specific processes of the company, thus failing to assess the actual quality of process safety.
To be truly effective, indicators must be clearly defined with respect to their scope of application: it must be stated upfront whether they concern process safety or personal safety, in order to properly evaluate the relevant prevention and risk management efforts. Although organizations collect a wealth of proactive data, they often lack the motivation or tools to make use of them, and it is often difficult to demonstrate the predictive power of such data. In this context, machine learning approaches can assist in identifying and even designing new leading indicators (Poh et al., 2018; Gondia et al., 2023), paving the way for more targeted and effective prevention.
@@ 2.2 Predictive models for occupational accident prevention
In recent years, the use of predictive models based on machine learning has become increasingly widespread in occupational safety, thanks to their ability to identify leading indicators (Gondia et al., 2023; Poh et al., 2018) and extract various risk factors (Kang and Ryu, 2019; Choi et al., 2020). Organizations collect vast amounts of data without always being able to detect the weak signals that would help initiate relevant preventive actions (Mearns, 2009; Tixier et al., 2016). To address this challenge, several algorithms have been deployed, including logistic regression, decision trees, random forests, boosting models, and neural networks (Kim et al., 2024). These techniques are applied across many sectors, such as construction (Tixier et al., 2016; Gondia et al., 2023; Poh et al., 2018), maritime transport (Kretschmann, 2020), metallurgy (Sarkar et al., 2020), and the service industry (Matías et al., 2008).
In the construction sector, Poh et al. (2018) compare various algorithms including logistic regression, decision trees, random forests, and SVMs to classify construction sites according to their safety level. The results show that random forests outperform the other models (see Table 8 in Poh et al. (2018)). Similarly, Gondia et al. (2023) use five predictors such as site environment, hazard exposure, human error, familiarity with the site, and current month to test algorithms such as naive Bayes, decision trees, random forests, SVMs, neural networks, and an ensemble model based on weighted voting. The ensemble approach yields better performance than any individual component (see Gondia et al., 2023, Table 7). The resulting prediction probabilities are used as leading indicators to assess site-level risk and enhance prevention efforts. Kretschmann (2020) also explores accident forecasting, introducing inspection-based indicators to anticipate safety conditions on ships, and using random forests for prediction.
National databases have also been used to identify the workers most exposed to fatal accidents. For example, Koc et al. (2023) analyzes 338,173 accidents in the Turkish construction sector using a combination of random forests, particle swarm optimization, and SHAP analysis, highlighting the importance of age, job position, experience, salary, and accident history. Similarly, Choi et al. (2020) leverage a large Korean dataset to predict fatality risks, comparing several models and confirming the superiority of random forests. These studies demonstrate that integrating national data and detailed worker-level indicators (e.g., age, role, seniority) enhances the ability to identify high-risk situations and key contributing factors.
In a recent study, Kim et al. (2024) explore deep learning approaches, including deep neural networks (DNN), long short-term memory (LSTM), and recurrent neural networks (RNN), to estimate the probability of fatalities in the context of natural disasters. The proposed model combines geographical and climatic data with conventional construction site information. After evaluating 36 different architectures, the authors identify a DNN optimized with the Adam algorithm as the most accurate (see Kim et al., 2024, Table 3).
@@ 2.3 Time series modeling and occupational accidents
A time series refers to a set of data collected at regular intervals, enabling the analysis of trends and the evolution of a phenomenon over time. In the context of occupational safety, such methods have primarily been applied at large scales over extended periods to uncover global trends, inform public policy, and compare the performances of companies in terms of accident prevention strategies (Carnero and Pedregal, 2010; Melchior et al., 2021).
Numerous studies rely on classical statistical models to investigate workplace accidents. Melchior et al. (2021) use various ARMA variants to estimate monthly mortality rates while Carnero and Pedregal (2010) and Verma et al. (2023) employ ARIMA and unobserved components models to forecast incident frequencies. Nazaripour et al. (2018) and Li et al. (2016) propose global indices designed to anticipate risk. Nazaripour et al. (2018) develop the customized predictive risk index (CPRI) using AR and MA models to assess safety performance in a steel plant, while Li et al. (2016) introduce an early warning system that combines multiple composite indices with a $GM(1,1)$ model. In these approaches, defining and interpreting thresholds requires substantial domain expertise to appropriately guide preventive actions.
Some studies have focused on leveraging machine learning models to forecast accident time series. Koc et al. (2022) apply wavelet decomposition to handle data non-stationarity and then use several algorithms, including artificial neural networks (ANN), support vector regression (SVR), and multivariate adaptive regression splines (MARS), to predict the daily number of accidents over short-, medium-, and long-term horizons. Their study relies on 393,160 construction-related accidents reported in Turkey between 2012 and 2020 and shows that integrating wavelets significantly improves forecast accuracy.

### v3

@@ Abstract
Introduction: Reducing the number of occupational accidents remains a major challenge for companies, as these events lead to significant human harm and financial losses. Although many organizations have implemented safety programs and made continuous efforts to improve their prevention strategies, these measures often remain insufficient to proactively and dynamically anticipate risks. In particular, safety inspections are still largely underexploited, and their integration into continuously updated predictive models has received little attention. Methods: we propose a model-agnostic framework for short-term occupational accident forecasting that leverages safety inspections and models accident occurrences as binary time series. The approach generates daily predictions, which are then aggregated into weekly safety assessments for better decision making. To ensure the reliability and operational applicability of the forecasts, we apply a sliding-window cross-validation procedure specifically designed for time series data, combined with an evaluation based on aggregated period-level metrics. Several machine learning algorithms, including logistic regression, tree-based models, and neural networks, are trained and systematically compared within this framework. Results: across all tested algorithms, the proposed framework reliably identifies upcoming high-risk periods and delivers robust period-level performance, demonstrating that converting safety inspections into binary time series yields actionable, short-term risk signals. Conclusions and Practical Applications: the proposed methodology converts routine safety inspection data into clear weekly and daily risk scores, detecting the periods when accidents are most likely to occur. Decision-makers can integrate these scores into their planning tools to classify inspection priorities, schedule targeted interventions, and funnel resources to the sites or shifts classified as highest risk, stepping in before incidents occur and getting the greatest return on safety investments.
@@ 1 Introduction
The international labour organization (ILO) estimates that nearly 300,000 people die each year due to occupational accidents (ILO2023). In France, the national health insurance recorded over 600 000 occupational accidents in 2023, nearly 700 of them fatal and this level has remained essentially flat for over a decade (Amelireports). This observation highlights the limitations of current strategies and the urgent need for new approaches to sustainably reduce both the frequency and severity of occupational accidents.
@@
Since Heinrich’s pioneering work (heinrich1931) and his domino theory, the understanding of occupational accidents has evolved considerably. These events are no longer seen as isolated or random, but rather as the outcome of a chain of contributing factors with complex interactions. Several theoretical models have emerged, such as Surry’s sequence of events model (surry1969), Reason’s Swiss cheese model (reason1990), and the cause tree method developed by the French INRS institute (inrs1970). While these models have helped structure accident investigation processes, they remain primarily retrospective, lack predictive capability, and often fail to capture the temporal dynamics of risk in complex environments (Qureshi2007; Larouzee2020).
@@
With the growing adoption of machine learning, new proactive strategies have been introduced in sectors such as construction, mining, agriculture, and services. Numerous studies demonstrate the potential of these techniques for incident prediction (SuarezSanchez2011; Rivas2011; Wang2018), risk assessment (Palei2009; Weng2011; Leu2013), injury severity classification (Chang2013; Esmaeili2015; Tixier2016), and risk-factor identification (Cheng2012; Amiri2016). However, most of these approaches rely on lagging indicator data collected after an accident, which limits their usefulness for anticipating risks and taking preventive action in advance. To address this limitation, the concept of safety leading indicators (SLIs) has gained momentum. In contrast to lagging indicators, SLIs allow for the early detection of weak signals based on proactive field data (Reiman2012; Hinze2013). Numerous studies emphasize their usefulness for constructing predictive models (Grabowski2007; Poh2018; Gondia2023). However, most of the existing frameworks using SLIs suffer from two major limitations: (1) they are rarely updated continuously, limiting their ability to adapt to evolving operational contexts; and (2) they often fail to explicitly capture temporal dependencies and to integrate recent information into the prediction process.
Time series models have also been applied to accident forecasting (carnero2010modelling; koc2022), but they typically operate at national or regional levels, over long time spans. Other works focus on building early warning systems based on composite indicators (LI201619; nazaripour2018safety), yet these systems require domain-expert thresholds and domain-specific calibration, limiting their operational flexibility. Despite these modeling advances, traditional methods remain predominant in occupational safety practice most notably the Fine–Kinney risk score (Fine1971; KinneyWiruth1976) and generic risk matrices (ISO31010). These approaches assign ordinal hazard indices from expert judgment and periodically aggregated indicators to prioritize controls. While simple and widely adopted, they depend on subjective scales and fixed thresholds and may compress quantitatively different risks (Cox2008). In this paper, we present a generic framework for short-term forecasting of occupational accidents at the company and department levels. We model daily accident occurrence as a binary time series (KedemFokianos2002; FokianosKedem2003). Unlike prior prediction and risk-scoring approaches, we explicitly learn day-to-day dynamics and cast the task as multi-output time-series classification. Conditioned on strictly ex-ante, continuously updated indicators from safety inspections, the model jointly predicts, for each site or department, the probability of at least one accident on each of the next $H$ days. We implement direct multi-horizon (MIMO) and direct-recursive (DirRec) strategies to produce probabilities for each forecast horizon. We convert these calibrated day-level probabilities into an interpretable weekly status: a day is flagged at risk when its probability exceeds a calibrated threshold, and a week is labeled risky if it contains at least one such day. This design leverages daily temporal dependence rather than aggregated counts, yields continuously updated risk maps across multiple horizons for planning instead of a single static score, and enables transparent, decision-oriented evaluation at every horizon. We assess performance with metrics suited to imbalanced classification (Luque2019) and validate robustness using sliding-window cross-validation tailored to time series (Tashman2000; bergmeir2012crossval; hyndman2018forecasting), which replicates the operational rollout procedure. For illustration, Table 1 shows the output of our approach applied to a specific department within the company under study, over a two-week period. The daily threshold for binarizing accident risk is set based on a calibration process and fixed here at $0.6$. In week 22, one day exceed this threshold, resulting in the classification of the entire week as risky. In contrast, in week 23, no day probability crosses the threshold, so the week is classified as safe. This example highlights how the approach can be used at the departmental level to provide timely and actionable insights for occupational risk management.
Week 22 Week 23
Day Mon Tue Wed Thu Fri Sat Sun Mon Tue Wed Thu Fri Sat Sun
Daily probability 0 0 0 0.65 0 0.39 0 0 0 0 0 0 0 0
Forecast (0/1) 1 1 0 0 0 0 0 0 0 0 0 0 0 0
@@
Table 1: Output of the proposed approach for a single department over two consecutive weeks. Weeks are flagged as risky or safe by comparing each day’s accident probability with a calibrated threshold of 0.6; week $2$ is labelled risky because two days exceed the threshold, whereas week 23 remains safe since no day does.
@@ 2 Literature review
An indicator is a qualitative or quantitative measure used to assess or monitor the evolution of a situation, phenomenon, or activity. In the field of data-driven occupational safety management, two main types of indicators are typically distinguished (Grabowski2007; Hopkins2009): lagging indicators and leading indicators. Lagging indicators such as accident or incident rates, compensation costs and number of injuries resulting in time off work (Choudhry2007; Hinze2013; Jazayeri2017) reflect the consequences of accidents that have already occurred. In other words, they are updated only after an accident happens. Several authors (Grabowski2007; Mengolini2008) argue that such indicators do not provide sufficiently useful information to prevent future accidents. According to Lindsay1992, a low number of reported accidents even over several years does not necessarily mean that risks are under control or that other incidents will not occur. Despite their limitations, these indicators remain widely used because they are easy to quantify and identify (Lingard2013; Almost2018) and allow organizations to benchmark against one another. (Elsebaei_2020).
In contrast, leading indicators provide early warning signs of accidents and adopt a more proactive approach, aiming to detect and act before incidents occur (Mearns2009; Eaton2013). Examples include near-miss report, safety talks, and safety inspections (Falahati2020). Field-level feedbacks can also be added to this list, as they enable quick and spontaneous collection of real-world operational data, helping to manage weak signals in real time. In the construction sector, Hinze2013 highlight the importance of these indicators and differentiates between passive leading indicators (e.g., number of employees trained or presence of a prevention plan) and active leading indicators (e.g., the percentage of safety meetings attended by supervisors). The latter reflects more accurately the dynamic reality of prevention efforts.
@@
Despite their promise, leading indicators remain difficult to adopt widely, partly due to the diversity of work environments: an effective indicator in construction may not apply in agriculture or maritime industries (Hinze2013; Xu2021). Furthermore, the subjectivity associated with some indicators such as the assessment of the severity of a hazardous situation can distort the actual perception of system performance or activity (Grabowski2007). Sometimes, the boundary between leading and lagging indicators is blurry, and some indicators are poorly defined or misaligned with their intended objectives. It is therefore crucial to distinguish between process safety hazards risks inherent to the operation of the system (e.g., explosions or toxic spills) and personal safety hazards which are more related to individual accidents such as falls, crushes, or electrocutions (Hopkins2009). Common lagging indicators like accident rates are often focused on personal safety and fail to capture process-related risks effectively. Similarly, some leading indicators (e.g., audit frequency) can remain too generic if they do not account for the specific processes of the company, thus failing to assess the actual quality of process safety.
To be truly effective, indicators must be clearly defined with respect to their scope of application: it must be stated upfront whether they concern process safety or personal safety, in order to properly evaluate the relevant prevention and risk management efforts. Although organizations collect a wealth of proactive data, they often lack the motivation or tools to make use of them, and it is often difficult to demonstrate the predictive power of such data. In this context, machine learning approaches can assist in identifying and even designing new leading indicators (Poh2018; Gondia2023), paving the way for more targeted and effective prevention.
2.2 Traditional safety scoring systems and manual risk assessments
Traditional safety scoring systems and manual risk assessments remain widely used in industry because they are simple, familiar, and inexpensive. In the Fine–Kinney method, risk is rated by multiplying ordinal scores for probability, exposure, and consequence to obtain a priority index for action (Fine1971; KinneyWiruth1976). Generic risk matrices, similarly map qualitative or semi-quantitative ratings to risk bands that guide decision making (ISO31010). These tools rely on expert judgment and indicators aggregated at periodic intervals (e.g., monthly audit findings or incident counts), which makes them practical for audits and routine reviews. However, they also have well documented limitations: scales are subjective and ordinal, fixed thresholds can distort prioritization, and combining categories may compress risks that differ in magnitude; more fundamentally, such schemes do not model day-to-day dynamics or provide horizon-specific accident probabilities, which are increasingly needed for proactive planning (Cox2008). In this study, we view these traditional methods as complementary: they remain useful for screening and communication, while our time series framework transforms leading indicators into daily, horizon-resolved probability forecasts that support short-term operational decisions.
@@ 2.3 Predictive models for occupational accident prevention
In recent years, the use of predictive models based on machine learning has become increasingly widespread in occupational safety, thanks to their ability to identify leading indicators (Gondia2023; Poh2018) and extract various risk factors (kang2019occupational_accidents; choi2020fatal_accidents). Organizations collect vast amounts of data without always being able to detect the weak signals that would help initiate relevant preventive actions (Mearns2009; Tixier2016). To address this challenge, several algorithms have been deployed, including logistic regression, decision trees, random forests, boosting models, and neural networks (kim2024deep). These techniques are applied across many sectors, such as construction (Tixier2016; Gondia2023; Poh2018), maritime transport (Kretschmann2020), metallurgy (Sarkar2020b), and the service industry (Matias2008).
In the construction sector, Poh2018 compare various algorithms including logistic regression, decision trees, random forests, and SVMs to classify construction sites according to their safety level. The results show that random forests outperform the other models (see Table 8 in Poh2018). Similarly, Gondia2023 use five predictors such as site environment, hazard exposure, human error, familiarity with the site, and current month to test algorithms such as naive Bayes, decision trees, random forests, SVMs, neural networks, and an ensemble model based on weighted voting. The ensemble approach yields better performance than any individual component (see Gondia2023, Table 7). The resulting prediction probabilities are used as leading indicators to assess site-level risk and enhance prevention efforts. Kretschmann2020 also explores accident forecasting, introducing inspection-based indicators to anticipate safety conditions on ships, and using random forests for prediction.
National databases have also been used to identify the workers most exposed to fatal accidents. For example, koc2023 analyzes 338,173 accidents in the Turkish construction sector using a combination of random forests, particle swarm optimization, and SHAP analysis, highlighting the importance of age, job position, experience, salary, and accident history. Similarly, choi2020fatal_accidents leverage a large Korean dataset to predict fatality risks, comparing several models and confirming the superiority of random forests. These studies demonstrate that integrating national data and detailed worker-level indicators (e.g., age, role, seniority) enhances the ability to identify high-risk situations and key contributing factors.
@@ 2.4 Time series modeling and occupational accidents
A time series refers to a set of data collected at regular intervals, enabling the analysis of trends and the evolution of a phenomenon over time. In the context of occupational safety, such methods have primarily been applied at large scales over extended periods to uncover global trends, inform public policy, and compare the performances of companies in terms of accident prevention strategies (carnero2010modelling; melchior2021forecasting).
Numerous studies rely on classical statistical models to investigate workplace accidents. melchior2021forecasting use various ARMA variants to estimate monthly mortality rates while carnero2010modelling and verma2023forecasting employ ARIMA and unobserved components models to forecast incident frequencies. nazaripour2018safety and LI201619 propose global indices designed to anticipate risk. nazaripour2018safety develop the customized predictive risk index (CPRI) using AR and MA models to assess safety performance in a steel plant, while LI201619 introduce an early warning system that combines multiple composite indices with a $GM(1,1)$ model. In these approaches, defining and interpreting thresholds requires substantial domain expertise to appropriately guide preventive actions.
Some studies have focused on leveraging machine learning models to forecast accident time series. koc2022 apply wavelet decomposition to handle data non-stationarity and then use several algorithms, including artificial neural networks (ANN), support vector regression (SVR), and multivariate adaptive regression splines (MARS), to predict the daily number of accidents over short-, medium-, and long-term horizons. Their study relies on 393,160 construction-related accidents reported in Turkey between 2012 and 2020 and shows that integrating wavelets significantly improves forecast accuracy.
@@
2.5 Deep learning for time series and occupational accident forecasting
Despite growing interest, deep learning remains comparatively underused in the occupational-accident literature; in particular, applying sequence models to binary day-level accident time series is still uncommon. Recent reviews indicate that accident prevention applications have focused mainly on computer vision and narrative text analysis rather than day-level forecasting of accident occurrence (Liu2022_AutCon). Nevertheless, there is accumulating evidence that modern neural approaches can add value in safety-related contexts. For example, kim2024deep explored deep neural networks (DNN), long short-term memory (LSTM), and recurrent neural networks (RNN) to estimate fatality probabilities under natural hazards by combining geographical, climatic, and construction site covariates. After benchmarking 36 architectures, they found that an Adam-optimized DNN attains the highest accuracy (kim2024deep, Table 3). Looking forward, several deep sequence-modeling families are directly applicable to day-level accident-risk forecasting. Thus temporal convolutional networks (TCN) use causal, dilated convolutions with residual connections to capture long effective memory with stable gradients and often challenge RNN or LSTM baselines on sequence benchmarks (Bai2018_TCN). Moreover, transformer-based time series models most notably the temporal fusion transformer (TFT) can produce multi-horizon forecasts alongside variable and horizon-level interpretability via gating, variable selection, and attention (Lim2021TFT).